# AI paintings vs. Human Paintings? Deciphering Public Interactions and Perceptions towards AI-Generated Paintings on TikTok

### v1

@@ Abstract
With the development of generative AI technology, a vast array of AI-generated paintings (AIGP) have gone viral on social media like TikTok. However, some negative news about AIGP has also emerged. For example, in 2022, numerous painters worldwide organized a large-scale anti-AI movement because of the infringement in generative AI model training. This event reflected a social issue that, with the development and application of generative AI, public feedback and feelings towards it may have been overlooked. Therefore, to investigate public interactions and perceptions towards AIGP on social media, we analyzed user engagement level and comment sentiment scores of AIGP using human painting videos as a baseline. In analyzing user engagement, we also considered the possible moderating effect of the aesthetic quality of Paintings. Utilizing topic modeling, we identified seven reasons, including looks too real, looks too scary, ambivalence, etc., leading to negative public perceptions of AIGP. Our work may provide instructive suggestions for future generative AI technology development and avoid potential crises in human-AI collaboration.
@@ Introduction
With the rise of various Generative AI technologies such as “Midjourney”, “stable diffusion” and “ChatGPT” in 2022, a vast array of AI-generated content (AIGC), such as images, texts, sounds, and even videos, has widely spread across major social media platforms. AIGC’s powerful functions and ease of use have gained the favor of many industries, particularly the art industry (Hitsuwari et al. 2023). Many famous artists, such as Sougwen Chung, Helena Sarin, Refik Anadol, etc., created advanced and groundbreaking artworks in cooperation with generative AI. As Sougwen Chung said [^1] https://zhuanlan.zhihu.com/p/149217204, “ As an artist working with these AI-generated tools, the prospect of artificial intelligence offers me a fresh perspective.” The AIGC has given the art industry a lot of new opportunities and provided artists with motivation.
However, the public has shown resistance towards genAI art. In December 2002, many painters uploaded images stating “NO TO AI-GENERATED IMAGES” [^2] https://arstechnica.com/information-technology/2022/12/artstation-artists-stage-mass-protest-against-ai-generated-artwork/ and called for a global boycott of AI-generated paintings (AIGP). This incident started because some artists felt that AI paintings had plagiarized their paintings without consent and their copyrights had been violated. This struggle lasted for nearly a year, and until now many painters are still calling for a ban on uploading AI-generated paintings on social media.
As AIGC circulates widely on social media, scholars have investigated how factors like text quality (Zhang and Gosline 2023) and AI usage (Messer 2024) affect viewers’ perceptions. However, many overlook AI’s painting capabilities and the aesthetic quality of AIGP on viewer participation and attitudes. Ragot et al. found significant cognitive bias in public perceptions of AIGP but did not consider Image Aesthetic Quality (IAQ) as a variable (Ragot, Martin, and Cojean 2020). Other studies (Bhandari, Chang, and Neben 2019; Shi, Huo, and Hou 2021) have shown that aesthetic quality significantly influences decision-making and perceptions. Therefore, we included IAQ as a potential moderating variable affecting user engagement.
@@
RQ1: How are people’s intentions to engage with AIGP on social media platforms?
@@
To answer these questions, we used human paintings as the baseline and compared the differences in user engagement levels and comment sentiments between AIGP and human paintings. To conduct the comparisons, we compiled a dataset with 207 AIGP and 210 human-painting image-text videos and over 80,000 comments on these videos from TikTok. Image-text video is a format similar to PPT, containing only images and text, where users swipe left or right to switch between images in the video. To answer RQ1, We used a formula to evaluate the user engagement level for every painting video. To answer RQ2, we adopted improved-aesthetic-predictor (LAION-Aesthetics V2333https://laion.ai/blog/laion-aesthetics/) as the image aesthetic quality assessment (IAQA) model to assess the IAQ of each painting for further comparison. To answer RQ3, we leveraged Natural Language Processing (NLP) techniques for sentiment analysis and topic modeling. The findings revealed that people still express stronger interaction intentions when browsing human paintings and that the IAQ does not significantly affect engagement intentions. We also found that people have more negative attitudes toward AIGP than human painting. The results of topic modeling listed seven main reasons for the negative sentiment. Early analysis of the public perception of AIGP will help the ICWSM research communities understand needs and articulate designs and interactive technologies that reduce plausible societal harm.
Figure 1: User interfaces of TikTok
@@
Social Attitudes toward Generative AI (Gen-AI)
@@
With the rapid development of AI, it has become increasingly important to study societal attitudes towards Gen-AI. Many researchers have also studied the attitudes of various groups of people towards generative AI technology in recent years. Grassini S et al. (Grassini 2023) developed and validated an Artificial Intelligence Attitude Scale (AIAS) designed to assess public perceptions of AI technology. Vasiljeva T et al. (Vasiljeva, Kreituss, and Lulle 2021)found that attitudes towards AI varied widely across industries. A significant difference in attitudes towards AI was found between employees of organizations implementing AI solutions and those without the intention of implementing AI solutions. Pinto et al.’s (Pinto dos Santos et al. 2019) survey of 263 medical undergraduates found that undergraduate medical students were not concerned about AI replacing human radiologists, and they understood the potential applications and impact of AI on radiology and medicine. YILDIZ T et al. (YILDIZ 2023)found that language learners were delighted and favored using AI in the learning process. All in all, people seem to have different attitudes towards AI technology based on different usage scenarios and different populations.
@@
AIGP, as the hot AI field of the moment, is developing at an astonishing rate, and the Paintings AI produces are already comparable to or surpass those created by humans. The technology not only brings us irreplaceable productivity and convenience (Xu 2019), but also the potential threat of indistinguishable deepfake problems (Bregler, Covell, and Slaney 2023), and Substitution of work  (Bendel 2023). Therefore, to determine the direction of future AI technological development and people’s needs, it is crucial and necessary to research the acceptance of artworks created by AI (Colton, Pease, and Saunders 2018). Many scholars have conducted social surveys on people’s attitudes towards AI-related work. Simon Colton (Colton 2008) was concerned about the perceived rejection of creative computers, but the evidence was unclear. Moffat and Kelly (Moffat and Kelly 2006) proved that significant bias for computer-generated music existed. Martin Ragot et al. (Ragot, Martin, and Cojean 2020) invited 565 participants to rate paintings created by generative AI tools and humans, respectively, and the results presented that the scores of human paintings were significantly higher than AI paintings. However, the Painting quality question in this study has not been solved, and the generative tools were incomplete. So, few studies constructed a complete process to analyze the attitude toward AIGP, primarily focusing on the painting artwork circulating on social media. In this research, we studied the social perception toward AIGP videos posted on social media and analyzed viewers’ sentiments from the interaction behaviors.
@@
Aesthetics, a significant branch of the visual arts, delves into exploring aesthetic categories such as beauty and its counterpart, ugliness (Zhu 2010). The aesthetic appeal of an image is evaluated based on universally accepted principles of photography. This appeal can be influenced by various elements, such as the strategic use of colors (Freeman 2007), the careful manipulation of contrast (Itten 1975), and the thoughtful arrangement of the image composition. Chatterjee and Leder et al. (Leder et al. 2004) introduced a five-stage, multi-level information processing model for image aesthetics. This model encompassed the perception of low-level information such as color, contrast, and complexity, the implicit integration of personal experiences and memories, explicit classification, cognitive mastery, and evaluation, culminating in aesthetic judgment and the generation of aesthetic emotions. However, computationally modeling this sequence for visual art images presents a significant challenge. Subsequently, Microsoft Asia Research Institute and Tsinghua University collaborated on a method to automatically differentiate photographs captured by professional photographers from those taken by customers (Tong et al. 2005). This study is widely recognized as the earliest research in the field of IAQA. The researchers employed a set of 21-class, 846-dimensional low-level global features to train a classification model for aesthetically categorizing the test images. Until 2014, with the emergence of AVA, a large-scale image aesthetic analysis dataset, image aesthetic analysis has been given an enormous boost based on deep learning technology (Jin et al. 2019). Lu X et al. (Lu et al. 2014) employed a dual-channel deep neural network to process local and global image blocks separately. Kong S et al. (Kong et al. 2016) developed the AADB aesthetic database, rated by human evaluators, and utilized deep CNN for aesthetic ranking and categorization of images. Chang KY et al. (Chan et al. 2019) created an aesthetic dataset, PCCD, and introduced a CNN and LSTM-based deep neural network capable of generating aesthetic linguistic comments on images. Jin X et al. (Jin et al. 2019) designed a novel deep convolutional neural network, ILGNet, for distinguishing between images of high and low aesthetic quality. Some famous datasets such as AVA (Aesthetic Visual Analysis) (Murray, Marchesotti, and Perronnin 2012), AADB (Aesthetics and Attributes Database ) (Kong et al. 2016) and CUHK-PQ (CUHK-Photo-Quality) (Ke, Tang, and Jing 2006) have collected billions of image samples. Therefore, using the aesthetic quality assessment models trained based on these datasets to evaluate the IAQ has been a good choice.
@@
We integrated several quantitative methods to solve and answer the proposed research questions. We compared user engagement and comment sentiment between 207 short videos related to AIGPs and 210 short videos showing human paintings. Additionally, to identify prominent reasons for negative comments, topic modeling was adopted to analyze the 42199 comments for AIGPs and 39763 comments for human paintings, respectively.
@@
Previous research has indicated that people generally hold negative attitudes towards AI-generated content, particularly in the realm of artworks. Therefore, our Hypotheses 1 and 3 are also based on this perspective, but we have narrowed the scope to AI-generated paintings (Wu et al. 2020). Meanwhile, visual aesthetic quality has been shown to be a significant factor in promoting user behaviors such as downloading and rating items on the Internet (Bhandari, Chang, and Neben 2019). Consequently, based on prior studies, we also posit that the aesthetic quality of AIGP positively impacts user engagement willingness (H2). To answer the research questions, we proposed the following three hypotheses:
@@
H2: AIGPs with high aesthetic quality scores lead to higher user engagement.
@@
Research Site
@@
TikTok is the fastest growing social network in the post-pandemic era, allowing users to publish and share short videos ranging from 15 to 60 seconds. It had 1506 million downloads in 2021 and has been the most downloaded social media application globally, surpassing Instagram, which had 1048 million downloads (Barta et al. 2023). In 2022, the average daily time spent on the TikTok mobile app by users worldwide was almost equivalent to that of Facebook and Instagram combined (Zhang et al. 2023). There are three main reasons for choosing TikTok as the survey platform for this study. First, as of December 15, 2023, videos under the tags related to AI painting on TikTok, such as “aiart”, “midjourny”, and “stable diffusion”, have received 10.7B, 7.5B, and 2.3B views. This not only indicates that a large amount of content related to AIGP has been widely disseminated on the TikTok platform, but also shows that users have a high degree of motivation and willingness to participate in discussions about AI-related topics. Second, based on TikTok’s recommendation algorithms, the platform analyzes users’ viewing preferences and directly pushes videos to them on the “For you” page (Zhao 2020). In this process, the viewer has no idea what the next video is about, thus avoiding the influence of users’ utilitarian participation on the video views in the experimental data. According to the Authentic Self Theory, individuals are more likely to exhibit genuine reactions when confronted with random events because they have limited time to prepare or control their responses (Vannini and Franzese 2008). Therefore, when a video with uncertain content like the one in TikTok suddenly “pops up” in front of the user, the user’s reactions are more likely to be genuine. Last but not least, according to TikTok’s platform management guidelines, all videos involving AI-generated content must be labeled, and videos without labels will not pass the platform’s review. Therefore, compared to other social media platforms such as Reddit, Instagram, and Twitter, TikTok is undoubtedly the most appropriate choice.
Data Collection and Preprocessing
Figure 2: Video Filtering Flowchart
@@
In our preliminary survey, we discovered that paintings on TikTok are presented in two formats: short videos and image-text videos. In short videos, many creators share their motivation for creating, record the creative process, or provide step-by-step painting tutorials. In these videos, user engagement is influenced by many factors unrelated to the artwork itself, such as the creator’s voice and behavior, the video’s length, etc. (Yang et al. 2022), which means that it is difficult to capture people’s actual attitudes towards the two types of paintings from these videos. Therefore, we selected image-text videos as our experimental objectives, which are similar to PPT format and contain only images and text. Users swipe left or right to switch between images in the video. In this kind of video, creators act more as sharers than narrators, only sharing their creations or AI-generated paintings. Thus, user engagement depends entirely on their understanding and feelings towards the artwork, unaffected by the creator’s value output. Therefore, choosing image-text videos for comparative analysis perfectly aligns with our research theme, making our experimental results more scientific and reliable.
Fig.2 a shows the initial steps for video collection and preprocessing. We created a new account on the TikTok video-sharing social media platform specifically for this study. By searching for tags related to AIGP and human paintings on TikTok, such as “Stable diffusion”, “Midjourney”, “AI paintings”, “Procreate” and “SAI” and scraping the web, we selected 2667 AI-labeled image-text videos and 1729 human painting image-text videos as of November 15, 2023. Considering videos with low views might not be representative, and some creators might restrict comment permissions, we excluded videos with fewer than 10k views, irrelevant content to the painting, and less than five comments. This refined our dataset to 778 AIGP videos and 935 human painting videos. To avoid favors or bias from users toward specific creators and repetitive Paintings, we adopted the same-author avoidance principle, ensuring no more than five works from the same author in our sample and excluding videos with repetitive Paintings within the Painting. Finally, after filtering based on these basic rules, we obtained 207 AIGP image-text videos and 210 human painting image-text videos.
Figure 3: Image Aesthetic Quality Schematic
@@
In this work, we define user engagement as the interactivity between users and videos (O’Brien and Toms 2008), reflecting the audience’s genuine interest and interaction with the videos. This is considered an indicator of the level of trust users have in a creator and their skill level. We adopted a formula for calculating TikTok user engagement on the HypeAuditor platform, an AI-driven analytical and discovery tool for Instagram, YouTube, TikTok, Twitter, and Twitch. It helps to find relevant influencers, understand their audience, increase the ROI of advertisers, and maintain the authenticity of influencer marketing. Many previous studies on social media data analysis have used this tool(Rogers ; Gil-Muñana and Sáez-Linero 2023). The formula for TikTok user engagement on Hype is $[(numberoflikes+numberofcomments+numberofshares)/numberofviews]X100\%$  [^4] https://hypeauditor.com/blog/what-is-tiktok-engagement-rate-why-brands-should-take-note/.

### v2

# AI vs. Human Paintings? Deciphering Public Interactions and Perceptions towards AI-Generated Paintings on TikTok

@@ Abstract
With the development of generative AI technology, a vast array of AI-generated paintings (AIGP) have gone viral on social media like TikTok. However, some negative news about AIGP has also emerged. For example, in 2022, numerous painters worldwide organized a large-scale anti-AI movement because of the infringement in generative AI model training. This event reflected a social issue that, with the development and application of generative AI, public feedback and feelings towards it may have been overlooked. Therefore, to investigate public interactions and perceptions towards AIGP on social media, we analyzed user engagement level and comment sentiment scores of AIGP using human painting videos as a baseline. In analyzing user engagement, we also considered the possible moderating effect of the aesthetic quality of Paintings. Utilizing topic modeling, we identified seven reasons, including hyperrealistic quality, ambivalent reactions, perceived theft of art, etc., leading to negative public perceptions of AIGP. Our work may provide instructive suggestions for future generative AI technology development and avoid potential crises in human-AI collaboration.
footnotetext: Corresponding author: Zhicong Lu, Email: zhiconlu@cityu.edu.hk.§§footnotetext: Current affiliation: School of Information Systems and Technology, University of New South Wales, Sydney, Australia. New Email: jiajun.wang5@student.unsw.edu.aufootnotetext: Current affiliation: University of Iowa, Iowa, United States.##footnotetext: Current affiliation: George Mason University, Virginia, United States. New Email: zlu6@gmu.edu
@@ Introduction
With the rise of various Generative AI technologies such as “Midjourney”, “stable diffusion” and “ChatGPT” in 2022, a vast array of AI-generated content (AIGC), such as images, texts, sounds, and even videos, has widely spread across major social media platforms. AIGC’s powerful functions and ease of use have gained the favor of many industries, particularly the art industry (Hitsuwari et al. 2023). Many famous artists, such as Sougwen Chung, Helena Sarin, Refik Anadol, etc., have created advanced and groundbreaking artworks in cooperation with generative AI. As Sougwen Chung said [^1] https://zhuanlan.zhihu.com/p/149217204, “ As an artist working with these AI-generated tools, the prospect of artificial intelligence offers me a fresh perspective.” The AIGC has given the art industry a lot of new opportunities and provided artists with motivation.
However, the public has shown resistance towards genAI art. In December 2022, many painters uploaded images stating “NO TO AI-GENERATED IMAGES” [^2] https://arstechnica.com/information-technology/2022/12/artstation-artists-stage-mass-protest-against-ai-generated-artwork/ and called for a global boycott of AI-generated paintings (AIGP). This incident started because some artists felt that AI paintings had plagiarized their paintings without consent and that their copyrights had been violated. This struggle lasted for nearly a year, and until now, many painters are still calling for a ban on uploading AI-generated paintings on social media.
As AIGC circulates widely on social media, scholars have investigated how factors like text quality (Zhang and Gosline 2023) and AI usage (Messer 2024) affect viewers’ perceptions. However, many overlook AI’s painting capabilities and the aesthetic quality of AIGP on viewer participation and attitudes. Ragot et al.  (Ragot, Martin, and Cojean 2020) found significant cognitive bias in public perceptions of AIGP, while Latikka et al.’s study  (Latikka et al. 2023) found that the public’s attitude toward AI in the field of art was generally negative. However, their research did not consider Image Aesthetic Quality (IAQ) as a variable. Other studies (Bhandari, Chang, and Neben 2019; Shi, Huo, and Hou 2021) have shown that aesthetic quality significantly influences decision-making and perceptions. Therefore, we included IAQ as a potential moderating variable affecting user engagement.
@@
RQ1: How are people’s engagement levels with AIGP on social media platforms?
RQ2: How does the image aesthetic quality (IAQ) influence people’s engagement intentions with AIGP and HP?
@@
To answer these questions, we used human paintings as the baseline and compared the differences in user engagement levels and comment sentiments between AIGP and human paintings. To conduct the comparisons, we compiled a dataset with 207 AIGP and 210 human-painting image-text videos and over 80,000 comments on these videos from TikTok. Image-text video is a format similar to PowerPoint, containing only images and text, where users swipe left or right to switch between images in the video. To answer RQ1, We used a formula to evaluate the user engagement level for every painting video. To answer RQ2, we adopted improved-aesthetic-predictor (LAION-Aesthetics V2333https://laion.ai/blog/laion-aesthetics/) as the image aesthetic quality assessment (IAQA) model to assess the IAQ of each painting for further regression. To answer RQ3, we leveraged Natural Language Processing (NLP) techniques for sentiment analysis and topic modeling. The findings revealed that people still express stronger interaction intentions when browsing human paintings and that the IAQ does not significantly affect engagement intentions. We also found that people have more negative attitudes toward AIGP than human painting. The results of topic modeling listed seven main reasons for the negative sentiment.
This paper makes contributions to the understanding of public perceptions towards AIGP on social media. Specifically, we contribute by:
  • Introducing image aesthetic quality (IAQ) as a critical moderating variable in the analysis of user engagement with AIGP and HP, thereby addressing a significant gap in the existing literature and providing novel insights into the dissemination patterns of AI-generated content on social media platforms.
  • Conducting a comprehensive and large-scale analysis of public perceptions towards AIGP in the context of social media, based on user engagement metrics, comment sentiment scores, and topic modeling, using human paintings as a baseline for comparison.
  • Identifying and categorizing seven key reasons behind negative public attitudes towards AIGP, including ‘Hyperrealistic Quality’, ‘Ambivalent Reactions’, ‘Failure to Meet Expectations’, and ‘Perceived Theft of Art’. These findings offer valuable guidance for the future optimization of AI generation tools to enhance user satisfaction and interaction experiences within human-computer interaction (HCI).
@@
Social Attitudes toward Generative AI (GenAI)
@@
With the rapid development of AI, it has become increasingly important to study societal attitudes toward GenAI. Many researchers have also studied the attitudes of various groups of people towards generative AI technology in recent years. Grassini S et al. (Grassini 2023) developed and validated an Artificial Intelligence Attitude Scale (AIAS) designed to assess public perceptions of AI technology. Vasiljeva T et al. (Vasiljeva, Kreituss, and Lulle 2021)found that attitudes toward AI varied widely across industries. A significant difference in attitudes toward AI was found between employees of organizations implementing AI solutions and those without the intention of implementing AI solutions. Pinto et al.’s (Pinto dos Santos et al. 2019) survey of 263 medical undergraduates found that undergraduate medical students were not concerned about AI replacing human radiologists, and they understood the potential applications and impact of AI on radiology and medicine. YILDIZ T et al. (YILDIZ 2023)found that language learners were delighted and favored using AI in the learning process. All in all, people seem to have different attitudes towards AI technology based on different usage scenarios and different populations.
@@
AIGP, as the hot AI field of the moment, is developing at an astonishing rate, and they are already comparable to or surpass those created by humans. The technology not only brings us irreplaceable productivity and convenience (Xu 2019) but also the potential threat of indistinguishable deepfake problems (Bregler, Covell, and Slaney 2023) and Substitution of work  (Bendel 2023). Therefore, to determine the direction of future AI technological development and people’s needs, it is crucial and necessary to research the acceptance of artworks created by AI (Colton, Pease, and Saunders 2018). Many scholars have conducted social surveys on people’s attitudes towards AI-related work. Simon Colton (Colton 2008) was concerned about the perceived rejection of creative computers, but the evidence was unclear. Moffat and Kelly (Moffat and Kelly 2006) proved that significant bias for computer-generated music existed. Martin Ragot et al. (Ragot, Martin, and Cojean 2020) invited 565 participants to rate paintings created by generative AI tools and humans, respectively, and the results presented that the scores of human paintings were significantly higher than AI paintings. However, the Painting quality question in this study has not been solved, and the generative tools were incomplete. So, few studies constructed a complete process to analyze the attitude toward AIGP, primarily focusing on the artwork circulating on social media. In this research, we studied the social perception toward AIGP videos posted on social media and analyzed viewers’ sentiments from the interaction behaviors.
Perception Differences between Human and AI Artwork
AI-driven images have demonstrated superior performance in producing a wide array of high-quality visuals, with some exhibiting richer color matching and more vivid detailing than human-created paintings. Given the similarity in visual effects, differences in people’s perceptions and attitudes toward these two types of artwork have become a significant topic of discussion. Existing literature has analyzed this issue from various perspectives. For example, eye-tracking studies have revealed an implicit negative bias toward AI-generated art compared to human-created art, suggesting that artistic creativity is still predominantly perceived as a human attribute (Zhou and Kawabata 2023). Across two field experiments, researchers found that participants preferred artworks labeled as “human-created” over those labeled as “AI-created,” even when all artworks were, in fact, generated by AI (Bellaiche et al. 2023). Another experiment similarly revealed that art labeled as human-made was rated higher on aesthetic dimensions, particularly when directly compared to AI-made art, thereby enhancing the perceived value of human effort (Horton Jr, White, and Iyengar 2023). To explore the underlying motivations behind these attitudes, one study showed that identical artworks were rated lower in terms of creativity and awe when labeled as AI-made rather than human-made. This labeling effect is one of the primary reasons for the reduced purchase intention toward AI-generated artwork (Millet et al. 2023).
Despite the valuable insights provided by existing studies, several gaps remain. First, most research has focused on face-to-face surveys and controlled experiments, with limited attention given to user perceptions and attitudes toward AI-generated art and human-created art within the specific context of social media platforms. This is important, as certain features of social media—such as user-driven dissemination (Kaplan and Haenlein 2010; Wang, Lai, and Kuan 2024) and the difficulty of verifying content accuracy (Pennycook and Rand 2019)—may significantly influence users’ decision-making processes. Second, in many studies, the crucial attribute of aesthetic quality has not been adequately controlled or considered.
@@
Aesthetics, a significant branch of the visual arts, delves into exploring aesthetic categories such as beauty and its counterpart, ugliness (Zhu 2010). The aesthetic appeal of an image is evaluated based on universally accepted principles of photography. This appeal can be influenced by various elements, such as the strategic use of colors (Freeman 2007), the careful manipulation of contrast (Itten 1975), and the thoughtful arrangement of the image composition. Chatterjee and Leder et al. (Leder et al. 2004) introduced a five-stage, multi-level information processing model for image aesthetics. This model encompassed the perception of low-level information such as color, contrast, and complexity, the implicit integration of personal experiences and memories, explicit classification, cognitive mastery, and evaluation, culminating in aesthetic judgment and the generation of aesthetic emotions. However, computationally modeling this sequence for visual art images presents a significant challenge. Subsequently, Microsoft Asia Research Institute and Tsinghua University collaborated on a method to automatically differentiate photographs captured by professional photographers from those taken by customers (Tong et al. 2005). This study is widely recognized as the earliest research in the field of IAQA. The researchers employed a set of 21-class, 846-dimensional, low-level global features to train a classification model for aesthetically categorizing the test images. Until 2014, with the emergence of AVA, a large-scale image aesthetic analysis dataset, image aesthetic analysis has been given an enormous boost based on deep learning technology (Jin et al. 2019). Lu X et al. (Lu et al. 2014) employed a dual-channel deep neural network to process local and global image blocks separately. Kong S et al. (Kong et al. 2016) developed the AADB aesthetic database, rated by human evaluators, and utilized deep CNN for aesthetic ranking and categorization of images. Chang KY et al. (Chan et al. 2019) created an aesthetic dataset, PCCD, and introduced a CNN and LSTM-based deep neural network capable of generating aesthetic linguistic comments on images. Jin X et al. (Jin et al. 2019) designed a novel deep convolutional neural network, ILGNet, for distinguishing between images of high and low aesthetic quality. Some famous datasets, such as AVA (Aesthetic Visual Analysis) (Murray, Marchesotti, and Perronnin 2012), AADB (Aesthetics and Attributes Database ) (Kong et al. 2016), and CUHK-PQ (CUHK-Photo-Quality) (Ke, Tang, and Jing 2006) have collected billions of image samples. Therefore, using the aesthetic quality assessment models trained based on these datasets to evaluate the IAQ has been a good choice.
@@
We integrated several quantitative methods to solve and answer the proposed research questions. We compared user engagement and comment sentiment between 207 image-text videos related to AIGP and 210 image-text videos showing human paintings. Additionally, to identify prominent reasons for negative comments, topic modeling was adopted to analyze the 42199 comments for AIGP and 39763 comments for human paintings, respectively.
Figure 1: Overview of the user interface elements in the video platform. A) The creator’s profile page on TikTok displays AIGI videos, showcasing thumbnails, view counts, and a label indicating that the video consists solely of images. B) The “For You” page highlights a selected video, presenting its associated metrics, including the number of likes, comments, favorites, and shares. C) The comments section contains AIGI-related discussions, showing the total number of comments, individual comment content, likes, and replies.
Figure 2: The data filtration process for AIGI and human painting videos. The left side (Step 1) illustrates the process of video selection and filtering based on content type, playing volume, and user engagement. The goal is to gather AIGI-related videos and ensure the dataset is clean and relevant for analysis. Videos are filtered using hashtags like “stable diffusion” or “AI-generated images,” and metadata such as view counts and upload dates are recorded. The right side (Step 2) demonstrates the aesthetic quality evaluation of extracted images and video categorization into four quality bins. This step focuses on assessing user interactions, such as likes, comments, and shares, to understand the popularity and reach of AIGI videos.
@@
Previous research has indicated that people generally hold negative attitudes toward AI-generated content, particularly in the realm of artwork (Hong and Curran 2019; Ragot, Martin, and Cojean 2020). Therefore, our Hypotheses 1 and 3 are also grounded in this perspective, although we have narrowed the scope specifically to AI-generated paintings (Wu et al. 2020). Meanwhile, visual aesthetic quality has been shown to be a significant factor in promoting user behaviors, such as downloading and rating items on the Internet (Bhandari, Chang, and Neben 2019).
In addition to the direct effect of IAQ on user engagement, we consider that IAQ may also play a moderating role in how users evaluate content from different sources, as suggested by existing literature. One study indicated that perceived aesthetics moderated trust and satisfaction in e-commerce environments (Cyr, Head, and Larios 2010). In the context of browser-based decision making, another study revealed that high aesthetic quality amplifies positive responses (Reinecke et al. 2013). Applying this mechanism to the context of our study, a higher level of IAQ may help offset users’ inherent bias against AI-generated artwork, leading viewers to develop greater acceptance of AI art and, consequently, encouraging interactive behaviors. To facilitate understanding of the moderating model in this study, we use “human origin level” to represent the creator type, with 0 defined as AIGP and 1 as HP. Therefore, we propose Hypothesis 3 that IAQ may negatively moderate the relationship between human origin level and user engagement. To answer these research questions, we proposed the following three hypotheses:
@@
H2: IAQ plays a negative moderating role in the relationship between human origin level and user engagement.
@@
Research Contexts
This study is based on user-generated content (UGC) from social media platforms. The choice of social media as the research context is primarily based on two main reasons. Firstly, social media provides real-time communication channels, enabling researchers to capture immediate and authentic interactions and perceptions from users. According to the Authentic Self Theory, individuals tend to express themselves more freely on social media, as they are less constrained by traditional societal expectations and norms. This relatively anonymous and disinhibited environment encourages users to present a more authentic self and to express their genuine attitudes (Vannini and Franzese 2008). Secondly, social media brings together users from diverse demographic and geographic backgrounds, allowing researchers to effectively analyze broad public opinions and perceptions (Perrin and Anderson 2019). This helps mitigate potential biases, such as regional differences. Moreover, social media offers access to large-scale datasets, allowing for more comprehensive and reliable analysis. Therefore, social media platforms are highly suitable as the research context for this study.
In addition, among many social media platforms, TikTok has become the fastest-growing social network in the post-pandemic era, allowing users to publish and share short videos ranging from 15 to 60 seconds. It recorded 1.506 billion downloads in 2021, surpassing Instagram, which had 1.048 billion downloads (Barta et al. 2023). In 2022, the average daily time users worldwide spent on the TikTok mobile app was nearly equivalent to the combined time spent on Facebook and Instagram (Zhang et al. 2023).
@@
There are three main reasons for choosing TikTok as the survey platform for this study. First, as of December 15, 2023, videos under tags related to AI painting on TikTok, such as “aiart”, “midjourney”, and “stable diffusion”, had accumulated 10.7B, 7.5B, and 2.3B views, respectively. This indicates that a large amount of AIGP-related content has been widely disseminated on TikTok, and that users show a high level of motivation and willingness to engage in discussions on AI-related topics. Second, TikTok’s recommendation algorithm analyzes users’ viewing preferences and pushes content directly to them on the “For You” page (Zhao 2020). In this process, users are unaware of what video will appear next, which reduces the influence of utilitarian behavior on video engagement metrics in the experimental data. According to the Authentic Self Theory, individuals are more likely to exhibit genuine reactions when confronted with random events, as they have limited time to prepare or control their responses (Vannini and Franzese 2008). Therefore, when a video with unpredictable content appears suddenly on TikTok, the user’s reaction is more likely to be authentic. Last but not least, according to TikTok’s platform management guidelines, all videos involving AI-generated content must be labeled. Videos without proper labeling will not pass the platform’s review. Fig. 1 shows an overview of TikTok’s user interface elements. Therefore, compared with other social media platforms such as Reddit, Instagram, and Twitter, TikTok is undoubtedly the most appropriate choice.
Data Collection and Pre-processing
@@
In our preliminary survey, we discovered that paintings on TikTok are presented in two formats: short videos and image-text videos. In short videos, many creators share their motivation for creating, record the creative process, or provide step-by-step painting tutorials. In these videos, user engagement is influenced by many factors unrelated to the artwork itself, such as the creator’s voice and behavior, the video’s length, etc. (Yang et al. 2022), which means that it is difficult to capture people’s actual attitudes toward the two types of paintings from these videos. Therefore, we selected image-text videos as our experimental objectives, which are similar to PowerPoint format and contain only images and text. Users swipe left or right to switch between images in the video. In this kind of video, creators act more as sharers than narrators, only sharing their creations or AI-generated paintings. Thus, user engagement depends mainly on their understanding and feelings towards the artwork, unaffected by the creator’s value output. Therefore, choosing image-text videos for comparative analysis perfectly aligns with our research theme, making our experimental results more scientific and reliable.
We created a new account on the TikTok video-sharing social media platform specifically for this study. By searching for tags related to AIGP and human paintings on TikTok, such as “Stable Diffusion”, “Midjourney”, “AI paintings”, “Procreate”, and “SAI”, and by scraping the web, we selected 2,667 AI-labeled image-text videos and 1,729 human painting image-text videos as of November 15, 2023. The selection of these search terms was guided by their direct relevance to the intersection of AI technology and digital art creation. They encompass both the AI tools used to generate art and the platforms where such art is further refined or utilized, ensuring a comprehensive exploration of public interactions and perceptions related to AI-generated paintings.
Considering that videos with low view counts might not be representative and that some creators may restrict comment permissions, we excluded videos with fewer than 10,000 views, videos irrelevant to painting, and those with fewer than five comments. There are two main reasons for this exclusion. First, videos with fewer than five comments do not provide sufficient data to support subsequent sentiment and semantic analyses. Second, video view counts on TikTok are largely influenced by the platform’s recommendation algorithm (Klug et al. 2021). The biases inherent in this algorithm may lead to a mismatch between the actual quality of a video and its level of exposure. For example, some high-quality videos may not receive adequate visibility, resulting in incomplete or unreliable user-related data. This, in turn, could interfere with the accuracy of our data analysis. Therefore, to minimize the potential impact of algorithmic mechanisms on the experimental results, such videos were excluded from the dataset. This filtering process refined our dataset to 778 AIGP videos and 935 human painting videos. To avoid user favoritism or bias toward specific creators and to prevent the inclusion of repetitive paintings, we adopted a same-author avoidance principle. Specifically, we limited the sample to no more than five works per author and excluded videos containing repetitive paintings within the same video.
Finally, after applying all the filtering criteria, we obtained a dataset comprising 207 AIGP image-text videos and 210 human painting image-text videos. The holistic data collection and filtration process is illustrated in Fig. 2.
Figure 3: Flowchart for calculating and comparing user engagement. This flowchart illustrates the process of calculating user engagement for AI-generated painting videos (N = 207) and human painting videos (N = 210). Engagement metrics, including number of views, number of comments, number of likes and number of shares, were normalized using RobustScaler and Min-Max scaling. The resulting user engagement scores were compared using the Independent T-test.
@@
In this work, we define user engagement as the interactivity between users and videos (O’Brien and Toms 2008), reflecting the audience’s genuine interest and interaction with the videos. We adopted a formula for calculating TikTok user engagement on the HypeAuditor platform, an AI-driven analytical and discovery tool for Instagram, YouTube, TikTok, Twitter, and Twitch. It helps to find relevant influencers, understand their audience, increase the ROI (Return On Investment) of advertisers, and maintain the authenticity of influencer marketing. Many previous studies on social media data analysis have used this tool(Trevisan et al. 2019; De Brito Silva et al. 2022). The formula for TikTok user engagement on Hype is $[(numberoflikes+numberofcomments+numberofshares)/numberofviews]X100\%$  [^4] https://hypeauditor.com/blog/what-is-tiktok-engagement-rate-why-brands-should-take-note/. Previous studies on TikTok engagement have also adopted these three metrics(Wang et al. 2024). The flowchart for calculating and comparing user engagement is shown in Fig. 3